OpenStudy (anonymous):

Tickets for the fall dance are $5 per person or $8 for couples. In order to cover the expenses, at least $1200 worth of tickets must be sold. Which is the correct inequality for this situation? a. 5c + 8s > 1200 c. 5s + 8c = 1200 b. 5c + 8s = 1200 d. 5s + 8c > 1200

OpenStudy (anonymous):

does the couples represent c and per person represent s ?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Write “Three times a number minus seven is less than 17 and greater than 5”, as a compound inequality. a. 5 < 3n - 7 < 17 c. 17 < 3n - 7 > 5 b. 3n-7 < 17 > 5 d. 3n-7 > 5 < 17

OpenStudy (anonymous):

5 < 3n - 7 < 17

OpenStudy (anonymous):

do u mind helping me for a little

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

yay thank you

OpenStudy (anonymous):

Jeff’s car averages 18 miles per gallon of gasoline. Select the greatest number of gallons of gasoline that he will need if he travels no more than 450 miles? a. 15 gallons c. 25 gallons b. 20 gallons d. 30 gallons

OpenStudy (anonymous):

450/18 = 25 gallons

OpenStudy (anonymous):

all real numbers a that are less than –6 or greater than 10 a. –6 < a < 10 c. a < 10 or a > –6 b. a < –6 or a > 10 d. –6 > a > 10

OpenStudy (anonymous):

-6 > a > 10

OpenStudy (anonymous):

–5x – 7 < 28

OpenStudy (anonymous):

-5x - 7 < 28 -5x < 28 + 7 -5x < 35 x > - 7

OpenStudy (anonymous):

Write an inequality and solve this problem. Two thirds of a number plus eight is greater than 12.What is the solution set? a. n > 6 c. n < 18 b. n > 18 d. n < 6

OpenStudy (anonymous):

2/3x + 8 > 12 2/3x > 12 - 8 2/3x > 4 x > 4/(2/3) x > 4 * 3/2 x > 12/2 x > 6

OpenStudy (anonymous):

If the farmer sells 210 boxes of oranges and 150 boxes of grapefruit, will he earn at least $4,000? If not, how many more boxes of oranges must he sell to make up the difference? a. yes c. no; 44 more boxes b. no; 65 more boxes d. no; 520 more boxes

OpenStudy (anonymous):

does it say how much the oranges and grapefruit are ?

OpenStudy (anonymous):

A citrus farm sells boxes of navel oranges for $8.00 and boxes of grapefruit for $12.00. The farmer wants to earn at least $4,000.

OpenStudy (anonymous):

210(8) + 150(12) = 1680 + 1800 = 3480....he will not make 4000 difference is 4000 - 3480 = 520 520/8 = 65....he would have to sell 65 more boxes of oranges to reach 4000

OpenStudy (anonymous):

Megan and Juan like to collect coins. Megan has 3 times as many coins as Juan. Megan buys 5 more coins. Juan has 3 coins. After Megan buys 5 more coins, how many more coins would Juan need to buy in order to have more coins than Megan? Let 'c' represent the number of coins that Juan needs to buy to have as many as Megan. a. c < 11 c. c > 11 b. c > 6 d. c > 5